40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog  Not logged in ELOG logo
Message ID: 185     Entry time: Fri May 16 00:13:36 2014
Author: Koji 
Type: Optics 
Category: Characterization 
Subject: Cavity mirror gluing part 1 

BS1/FM1/FM2 for I1OMC were glued.

FM1 had to be intentionally rotated.
FM1 had to be intentionally shifted to avoid scattering spot.

Pin: 36.3 / Ptrans: 33.7 = Raw transmission 92.8%
Vunlock = 6.30 / Vlock = 0.120

Mode matching (estim) 0.98
Loss per mirror 84ppm
Cavity transmission 0.947

ummm

Tomorrow:
- Transmission needs to be optimized
- Apply 50V to a PZT
- Cavity FSR/HOM should be optimized
- gluing

Put a cover
Return power meter / DC supply

ELOG V3.1.3-